OpenStudy (anonymous):

What are the square roots of 256? a 13 and –13 b 15 and –15 c 16 and –16 d 128 and –128

OpenStudy (anonymous):

use calculator

OpenStudy (anonymous):

i did i only get 1 answer!

OpenStudy (pigrenex):

C

OpenStudy (anonymous):

calculator wont give the answer holding negative sign

OpenStudy (anonymous):

@pigrenex how do you get that?

OpenStudy (anonymous):

a negative squared will always be positive

OpenStudy (maheshmeghwal9):

But in real maths there are 2 answers +16 & -16 since when both squared , we get 256

jimthompson5910 (jim_thompson5910):

Notice how 16*16 = 256 and (-16)*(-16) = 256 So this points us to the answer of C

OpenStudy (maheshmeghwal9):

hence , C

OpenStudy (anonymous):

but whenever you will try to find out the squareroots of a number you will get two..

OpenStudy (pigrenex):

-16 x -16 = 256 every time you multiply equal signs, you get a positive answer

OpenStudy (maheshmeghwal9):

(+)*(+)=(+) & (-)*(-)=(+)

OpenStudy (anonymous):

ok i got it thanks every one :)

OpenStudy (lgbasallote):

a fun way to solve \(x^2 = 256\) \(x^2 - 256 = 0\) \((x+16)(x-16) = 0\) \(x = 16; x = -16\)

OpenStudy (maheshmeghwal9):

wc:)